given sin theta =-1/2 and sec theta >0,find the exact value of tan 2theta
Answer by lwsshak3(6472)

given sin theta =-1/2 and sec theta >0,find the exact value of tan 2theta
**
sinx<0 and secx>0 only in quadrant IV
In quadrant IV:
tanx=-1/√3
Identity:
tan(2x)=(2tanx)/(1-tan^2x)
=(-2/√3)/1-(1/3)
=-2√3/2/3
=-6/2√3
=-√3*6/6
=√3